OpenStudy (anonymous):

tan^2 (x)=(1-cosx)/(1+cosx) true or false

OpenStudy (anonymous):

true?

OpenStudy (primeralph):

Multiply top and bottom by 1-cosx.

OpenStudy (anonymous):

\[\frac{ 1-cosx }{ 1+cosx } * \frac{1-cosx}{1-cosx} = \frac{1-2cosx + \cos^{2}x}{1-\cos^{2}x}\]

OpenStudy (anonymous):

now solve the rest with substitution and factor.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got some weird thing when i factor it

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

you should get cosx-1 and cosx -1

OpenStudy (anonymous):

never mind i figured it out